The Bell Weather Co. is a new firm in a rapidly growing industry. The company is planning on increasing its annual dividend by 15 percent a year for the next 4 years and then decreasing the growth rate to 6 percent per year. The company just paid its annual dividend in the amount of $2.90 per share. What is the current value of one share of this stock if the required rate of return is 8.40 percent?

Multiple Choice

A. $175.72

B. $162.24

C. $226.92

D. $178.62

E. $224.02
Answer #1

D1=(2.9*1.15)=3.335

D2=(3.335*1.15)=3.83525

D3=(3.83525*1.15)=4.4105375

D4=(4.4105375*1.15)=5.07211813

Value after year 4=(D4*Growth rate)/(Required rate-Growth rate)

=(5.07211813*1.06)/(0.084-0.06)

=224.018551

Hence current value=Future dividend and value*Present value of discounting factor(rate%,time period)

=3.335/1.084+3.83525/1.084^2+4.4105375/1.084^3+5.07211813/1.084^4+224.018551/1.084^4

=$175.72 per unit.